OpenStudy (anonymous):

3c + 15 --- --- c-5 5-c

OpenStudy (anonymous):

Factor out a -1 and your denominators will be the same

OpenStudy (anonymous):

3c/c-5 - 15/c-5 = 3c-15 /c-5 now factor to simplest form

OpenStudy (anonymous):

3(c-5)/c-5 = 3

OpenStudy (anonymous):

I factored out the greatest common factor of 3c-15, which was 3.

OpenStudy (anonymous):

oh right thanks so much!

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4x 2 --- - ---- (x+3)^2 x+3

OpenStudy (anonymous):

The common denominator will be (x+3)(x+3)

OpenStudy (anonymous):

4x + 2(x+3) is your numerator.

OpenStudy (anonymous):

its 4x-2

OpenStudy (anonymous):

4x+2x+6 = 6x+6 factor out 6 and you get 6(x+1)/(x+3)^2

OpenStudy (anonymous):

oops...didn't see subtraction

OpenStudy (anonymous):

4x-2x-6 = 2x-6 = 2(x-3) for numerator

OpenStudy (anonymous):

wouldnt it be 4x - 2(x+3)

OpenStudy (anonymous):

Yes, but I distributed the -2 into the parenthesis

OpenStudy (anonymous):

That gives the 4x-2x-6

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

2(x-3) _______ (x+3)^2

OpenStudy (anonymous):

thanks so much that helped a lot!!!

OpenStudy (anonymous):

You're welcome